2. Calculate the force of elasticity (or tension) with which an 8 kg hanging candlestick pulls the cord. How much is the deformation of the cord when it is known that the coefficient of elasticity has values ​​k = 4 × 10⁴ N / m.


1
Expert's answer
2021-01-14T10:39:36-0500

(a) We can find the force of tension from the Newton's Second Law of Motion:


"\sum F_y=ma_y=0,""T-mg=0,""T=mg=8\ kg\cdot 9.8\ \dfrac{m}{s^2}=78.4\ N."

(b) Finally, we can find the deformation of the cord from the Hooke's Law:


"T=F=kx,""x=\dfrac{T}{k}=\dfrac{78.4\ N}{4\cdot10^4\ \dfrac{N}{m}}=19.6\cdot10^{-4}\ m."

Answer:

(a) "T=78.4\ N."

(b) "x=19.6\cdot10^{-4}\ m."